OpenStudy (samsan9):

what is the length of the major axis of the ellipse shown below? (x-1)^2/9+(y+5)^2/16=1

OpenStudy (samsan9):

\[\frac{ (x-1)^2 }{ 9 }+\frac{ (y+5)^2 }{ 16 }=1\]

OpenStudy (samsan9):

A.) 16 B.) 8 C.) 32 D.) 4 i believe its 4 but i wanted to be sure

OpenStudy (samsan9):

or maybe it is B